OpenStudy (anonymous):

Write the equation of a line in point-slope form that has a slope of and passes through the point (2, -7).

OpenStudy (anonymous):

what's the slope?

OpenStudy (anonymous):

it dont give me a slope

OpenStudy (anonymous):

dy/dx gives the slope: so if the points are (2,-7) and (x,y) then the slope is given by: y- (-7)/ x-2 =(y+7)/(x-2) to get the equation you have to equate this to a given slope which unfortunately you weren't provided with

OpenStudy (anonymous):

sorry the slope is 2/3

OpenStudy (anonymous):

then just equate (y+7)/(x-2) to 2/3

OpenStudy (anonymous):

u aint makin sense

OpenStudy (anonymous):

OK step by step: (y+7)/(x-2) =2/3 cross multiply; 2(x-2) = 3(y+7) 2x-4 =3y+21 2x-3y = 25 remember equation of a line is given as y =mx +c where m is the gradient and c the intercept so 2x-3y = 25 becomes -3y = 25 - 2x 3y = 2x -25 y = 2/3x - 25/8 is that clearer?

OpenStudy (anonymous):

thank you thank you thank you now i understand pedro

OpenStudy (anonymous):

welcome